TIME-RESOLVED ENERGY TRANSFER STUDIES OF LOW TEMPERATURE PARA-NH_{3} - He COLLI-SIONS

\underline{D. R. WILLEY}, R. E. TIMLIN, JR., M. C. DEAN, Z. S. CATALADI, Department of Physics, Allegheny College, Meadville, PA 16335.

We report results of low temperature (5-15 K) experimental studies of para-NH_{3} in collisions with He using time-resolved MW - MW techniques. In these studies, ammonia is collisionally cooled in a low temperature cell filled with a background of helium. A short pump pulse is used to excite a (J, K ) inversion pair while a second weak source probes nearby inversion transitions. For para-NH_{3} only the (1, 1) and (2, 2) levels have significant population at these temperatures. Adding in the inversion doublets, this is effectively a four-level system, making it feasible to extract approximate state-to-state collision rates. We will present our experimental results and compare them to state-to-state cross sections calculated using recent NH_{3}-He potential surfaces.